# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 3.3.3 Field Equations for $g^{mn}$
where the equality between the definition of $V^{a}$ and the expression its second line is proved in appendix C.1. As proved in appendix C.2, a variation with respect to the Levi-Civita connection can be expressed as a variation with respect to the metric as
$\sqrt{g}A_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\sqrt{g}\tfrac{1}{2}\left[\hat{\nabla}_{a}(A_{(mn)}{}^{a}-A^{a}{}_{mn})\right]\delta g^{mn}$ (3.65)
$\sqrt{g}A^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $\sqrt{g}\tfrac{1}{2}g_{mn}(\bar{\nabla}_{a}A^{a})\delta g^{mn}$ (3.66)
up to total derivatives and where $A_{c}{}^{ab}=A_{c}{}^{ba}$ is an arbitrary tensor symmetric in its last two indices and $A^{a}$ is an arbitrary vector. Using this, Eq. (3.62) becomes
@@
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\left[g_{mn}{\hat{\nabla}}_{a}V^{a}+{\hat{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\right]\delta g^{mn}~{}~{}~{}.$ (3.67)
Using Eqs. (3.2)  and (3.8) to switch to ¯${\bar{\nabla}}_{a}$ we have
@@
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\Big{[}g_{mn}{\bar{\nabla}}_{a}V^{a}+{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})$ (3.68)
$+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\Big{]}\delta g^{mn}~{}~{}~{}.$ (3.69)
@@
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (3.70)
@@
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $J_{0}\Big{[}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\tfrac{1}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})$ (3.71)
$\hskip 50.0pt+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\Big{]}\delta g^{mn}~{}~{}~{}.$ (3.72)
The rest of the terms involving $\delta g^{mn}$ will be found by expanding the action, Eq. (3.31), to expose the factors of the inverse metric $g^{mn}$
$\mathcal{L}=$ $-\frac{1}{2\kappa_{0}}g^{mn}\mathcal{K}_{mn}+J_{0}\lambda_{0}^{2}g^{ab}g^{cm}g^{dn}K_{bcd}K_{amn}$ (3.73)
$-J_{0}\left(g_{am}g^{bn}g^{cp}g^{dq}\mathcal{K}^{a}_{\ bcd}\mathcal{K}^{m}{}_{npq}-4g^{am}g^{bn}\mathcal{K}_{ab}\mathcal{K}_{mn}+g^{ab}g^{mn}\mathcal{K}_{ab}{\mathcal{K}}_{mn}\right).$ (3.74)
Next, the variation begins by varying the product of the Lagrangian density with $\sqrt{g}$
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2}g^{mn}\delta g_{mn}\mathcal{L}-\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}{\mathcal{K}}_{mn}$ (3.75)
$+\sqrt{g}J_{0}\lambda_{0}^{2}(\delta g^{ab}K_{acd}K_{b}{}^{cd}+2\delta g^{dn}K_{bcd}K^{bc}{}_{n})$ (3.76)
$-\sqrt{g}J_{0}(\delta g_{am}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}^{mbcd}+\delta g^{bn}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}_{an}{}^{cd}+2\delta g^{dq}{\mathcal{K}}^{a}{}_{bcd}{\mathcal{K}}_{a}{}^{bc}{}_{q})$ (3.77)
$+\sqrt{g}J_{0}(8\delta g^{am}{\mathcal{K}}_{ab}{\mathcal{K}}_{m}{}^{b}+2\delta g^{mn}{\mathcal{K}}_{mn}{\mathcal{K}})+\sqrt{g}\delta_{\hat{\Gamma}}\mathcal{L}~{}~{}~{}.$ (3.78)
Recall here on the manifold $\mathcal{M}$, we raise and lower indices with $g^{mn}$ and $g_{mn}$ as for example ${\mathcal{K}}_{a}{}^{bc}{}_{q}=g_{am}g^{bn}g^{cp}{\mathcal{K}}^{m}{}_{npq}$. Using Eq. (3.71), collecting like terms, and relabeling indices we have
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}\bigg{\{}-{\mathcal{K}}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (3.79)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}\left(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n}\right)+2\kappa_{0}J_{0}\lambda_{0}^{2}\left(8{\mathcal{K}}_{mb}{\mathcal{K}}_{n}{}^{b}-2{\mathcal{K}}{\mathcal{K}}_{mn}\right)$ (3.80)
$+2\kappa_{0}J_{0}\left({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{a}{}_{bcn}\right)+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (3.81)
$+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}\right.$ (3.82)
$\hskip 100.0pt\left.+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]\bigg{\}}~{}~{}~{}.$ (3.83)
We collect the factor in the curly brackets in an equation of motion tensor $\mathbb{H}_{mn}$
$\delta_{g}(\sqrt{g}\mathcal{L})=$ $\tfrac{\sqrt{g}}{2\kappa_{0}}\delta g^{mn}\mathbb{H}_{mn}$ (3.84)
@@
$0=\mathbb{H}_{mn}=$ $\mathcal{H}_{mn}+\tfrac{2\kappa_{0}J_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+2\kappa_{0}J_{0}\left[{\bar{\nabla}}_{a}(\tilde{\mathbb{E}}_{(mn)}{}^{a}-\tilde{\mathbb{E}}^{a}{}_{mn})\right.$ (3.85)
$\left.\hskip 120.0pt+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{|p|n)}{}^{a}+{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}_{n)p}{}^{a}-{\tilde{C}}^{p}{}_{a(m}\tilde{\mathbb{E}}^{a}{}_{n)p}\right]$ (3.86)
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}-\kappa_{0}g_{mn}\left[\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)+\mathcal{L}\right]$ (3.87)
$+2\kappa_{0}J_{0}\lambda_{0}^{2}(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n})+2\kappa_{0}J_{0}(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}{}_{n}-2{\mathcal{K}}{\mathcal{K}}_{mn})$ (3.88)
$+2\kappa_{0}J_{0}({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{abcn})~{}~{}~{}.$ (3.89)
@@
To show that $\mathcal{H}_{mn}$ is equivalent to the form in Eq. (2.65), Eq. (3.4) is used to relate the ¯${\bar{\nabla}}_{a}$ term to a ˘$\breve{\nabla}_{a}$ term as
¯${\bar{\nabla}}_{a}({\mathcal{K}}^{a}{}_{bcd}K^{bcd})=(\breve{\nabla}_{a}-({\rm d}+1)g_{a})({\mathcal{K}}^{a}{}_{bcd}K^{bcd})=(\nabla_{a}-({\rm d}+1)\text{\textcent}_{a})({\mathcal{K}}^{a}{}_{bcd}K^{bcd}).$ (3.90)
4 Classical Equivalence of Field Equations Across Formulations

### v3

@@ III.3.2 Field Equations for ${\tilde{C}}^{c}{}_{ab}$
where the equality between the definition of $V^{a}$ and the expression its second line is proved in appendix C.1. Using Eq. (108) we split the variation Eq. (170) into two parts
$\delta_{{\bar{\Gamma}}}\mathcal{L}=\delta_{{\hat{\Gamma}}}\mathcal{L}+\delta_{{\tilde{C}}}\mathcal{L}$ (173)
$\delta_{{\hat{\Gamma}}}\mathcal{L}=2J_{0}(\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\hat{\Gamma}}^{c}{}_{ab}+V^{a}\delta{\hat{\Gamma}}^{m}{}_{am})$ (174)
$\delta_{{\tilde{C}}}\mathcal{L}=2J_{0}\tilde{\mathbb{E}}_{c}{}^{ab}\delta{\tilde{C}}^{c}{}_{ab}$ (175)
where we have exploited the tracelessness of ${\tilde{C}}^{c}{}_{ab}$ to simplify $\delta_{\tilde{C}}\mathcal{L}$ to a single term involving $\tilde{\mathbb{E}}_{c}{}^{ab}$. Thus Eq. (175) yields the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and Eq. (174) leads to a contribution to the metric $g^{mn}$ field equations. Summarizing, the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are
$0=\tilde{\mathbb{E}}_{c}{}^{ab}\equiv$ $\mathbb{E}_{c}{}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}\mathbb{E}_{m}{}^{b)m}$ (176)
$\mathbb{E}_{c}^{\ ab}=$${\bar{\nabla}}_{c}\mathbb{K}^{ab}-{\bar{\nabla}}_{d}{\mathcal{K}}_{c}^{\ (ab)d}-\lambda_{0}^{2}{\bar{{\mathcal{P}}}}_{dc}K^{(ab)d}$ (177)
where recall the projectively invariant tensor $\mathbb{K}^{ab}$ was defined in Eq. (82). The equations of motion for ${\tilde{C}}^{c}{}_{ab}$ are covariant and manifestly projective invariant as they are composed entirely of such fields. Where the diffeomorphism field equations for ${\mathcal{D}}_{ab}$ are identical to those for ${\bar{{\mathcal{P}}}}_{ab}$, the equations of motion for $\Pi^{c}{}_{ab}$ are different from those for ${\tilde{C}}^{c}{}_{ab}$. This is clearly seen in Eq. (92) in that $\tilde{\mathbb{E}}_{c}^{\ ab}$ and ~$\tilde{\mathcal{E}}_{c}^{\ ab}$ differ by terms involving the diffeomorphism field equations tensor $\mathbb{D}^{ab}$. We rewrite this below in terms of $\tilde{\mathbb{E}}_{c}^{\ ab}$ vanishing instead of ~$\tilde{\mathcal{E}}_{c}^{\ ab}$, as this enforces the equations of motion for $\tilde{C}^{c}_{\ ab}$
$0=\tilde{\mathbb{E}}_{c}{}^{ab}=$$\tilde{\mathcal{E}}_{c}{}^{ab}-\left(g_{c}\mathbb{D}^{ab}-\tfrac{1}{{\rm d}+1}\delta_{c}{}^{(a}g_{m}\mathbb{D}^{b)m}\right)~{}.$ (178)
We see here that upon enforcing the diffeomorphism field equations (going on-shell)
$\mathbb{D}^{ab}=0\Rightarrow\tilde{\mathbb{E}}_{c}^{\ ab}=\tilde{\mathcal{E}}_{c}^{\ ab}=0$ (179)
thus the equations of motion for ${\tilde{C}}^{c}{}_{ab}$ and $\Pi^{a}{}_{ab}$ are identical on-shell.
III.3.3 Field Equations for $g^{mn}$
To derive the field equations for the metric, we must vary the product of $\sqrt{|g|}$ and the Lagrangian, Eq. (54)
$\delta_{g}(\sqrt{|g|}\mathcal{L})=$ $\tfrac{\sqrt{|g|}}{2}g^{mn}\delta g_{mn}\mathcal{L}-\tfrac{\sqrt{|g|}}{2\kappa_{0}}\delta g^{mn}{\mathcal{K}}_{mn}$ (180)
$+\sqrt{|g|}\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}+\sqrt{|g|}\delta_{\hat{\Gamma}}\mathcal{L}$ (181)
$=$ $\tfrac{\sqrt{|g|}}{2\kappa_{0}}\delta g^{mn}(-{\mathcal{K}}_{mn}-\kappa_{0}g_{mn}\mathcal{L})$ (182)
$+\sqrt{|g|}\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}+\sqrt{|g|}\delta_{\hat{\Gamma}}\mathcal{L}$ (183)
where in going from the first to the second equality, we have used the relationship
$\delta g_{mn}=-g_{ma}g_{nb}\delta g^{ab}~{}.$ (184)
The term involving $\delta_{g}\mathbb{L}$ is evaluated setting variations of the Levi-Civita connection $\delta\hat{\Gamma}^{c}_{\ ab}$ to zero to avoid double counting of such variations present in the last term involving $\delta_{\hat{\Gamma}}\mathcal{L}$.
We next focus on this last term, continuing with Eq. (174) working to rewrite it in terms of a variation with respect to $g^{mn}$. As proved in appendix C.2, a variation with respect to the Levi-Civita connection can be expressed as a variation with respect to the metric as
$\tilde{\mathbb{E}}_{c}{}^{ab}\delta\hat{\Gamma}^{c}{}_{ab}=$ $\tfrac{1}{2}\delta g^{mn}\hat{\nabla}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}$ (185)
$\tilde{\mathbb{E}}^{*\ \ a}_{mn}\equiv$ $\tilde{\mathbb{E}}_{(mn)}^{\ \ \ \ \ a}-\tilde{\mathbb{E}}^{a}_{\ mn}$ (186)
$V^{a}\delta\hat{\Gamma}^{m}{}_{am}=$ $\tfrac{1}{2}\delta g^{mn}g_{mn}\bar{\nabla}_{a}V^{a}$ (187)
up to total derivatives. Using this, Eq. (174) becomes
@@ III.3.3 Field Equations for $g^{mn}$
$\delta_{{\hat{\Gamma}}}\mathcal{L}=$ $\delta g^{mn}J_{0}\left(g_{mn}{\bar{\nabla}}_{a}V^{a}+{\hat{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}\right)~{}.$ (188)
Using Eqs. (105)  and (108) to switch to ¯${\bar{\nabla}}_{a}$ we have
@@
$\delta_{{\hat{\Gamma}}}\mathcal{L}$ $=\delta g^{mn}J_{0}\left[g_{mn}{\bar{\nabla}}_{a}V^{a}+{\bar{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}+{\tilde{C}}^{p}_{\ a(m}\tilde{\mathbb{E}}^{*\ \ a}_{n)p}\right].$ (189)
@@
¯${\bar{\nabla}}_{a}V^{a}=\tfrac{1}{{\rm d}+1}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (190)
@@
$\delta_{{\hat{\Gamma}}}\mathcal{L}$ $=\delta g^{mn}J_{0}\Big{[}-\tfrac{\lambda_{0}^{2}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (191)
$+\tfrac{1}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+{\bar{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}+{\tilde{C}}^{p}_{\ a(m}\tilde{\mathbb{E}}^{*\ \ a}_{n)p}\Big{]}.$ (192)
Our final task is to expand the third term of Eq. (180), $\sqrt{|g|}\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}$ We start by expanding the PGB Lagrangian itself, Eq. (55), to expose the factors of the metric $g_{mn}$ and its $g^{mn}$
$\mathbb{L}=$ $J_{0}(\lambda_{0}^{2}g^{ab}g^{cm}g^{dn}K_{bcd}K_{amn}-g_{am}g^{bn}g^{cp}g^{dq}\mathcal{K}^{a}_{\ bcd}\mathcal{K}^{m}_{\ npq})$
$+J_{0}(4g^{am}g^{bn}\mathcal{K}_{ab}\mathcal{K}_{mn}-g^{ab}g^{mn}\mathcal{K}_{ab}{\mathcal{K}}_{mn}).$ (193)
Variation of $\mathbb{L}$ with respect to the metric (but not Levi-Civita connection to prevent the aforementioned double counting) proceeds via several product rules resulting in
$\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}=J_{0}\lambda_{0}^{2}(\delta g^{ab}K_{acd}K_{b}^{\ cd}+2\delta g^{dn}K_{bcd}K^{bc}_{\ \ n})$
$-J_{0}(\delta g_{am}{\mathcal{K}}^{a}_{\ bcd}{\mathcal{K}}^{mbcd}+\delta g^{bn}{\mathcal{K}}^{a}_{\ bcd}{\mathcal{K}}_{an}^{\ \ cd}+2\delta g^{dq}{\mathcal{K}}^{a}_{\ bcd}{\mathcal{K}}_{a}^{\ bc}{}_{q})$
$+J_{0}(8\delta g^{an}{\mathcal{K}}_{ab}{\mathcal{K}}^{b}_{\ n}-2\delta g^{mn}{\mathcal{K}}_{mn}{\mathcal{K}})~{}.$ (194)
Recall here on the manifold $\mathcal{M}$, we raise and lower indices with $g^{mn}$ and $g_{mn}$ as for example ${\mathcal{K}}_{a}{}^{bc}{}_{q}=g_{am}g^{bn}g^{cp}{\mathcal{K}}^{m}{}_{npq}$. Using Eq. (184), collecting like terms, and relabeling indices we have
$\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}=\delta g^{mn}\left[J_{0}\lambda_{0}^{2}\left(K_{mcd}K_{n}{}^{cd}+2K_{bcm}K^{bc}{}_{n}\right)\right.$
$\quad+J_{0}\left({\mathcal{K}}_{mbcd}{\mathcal{K}}_{n}{}^{bcd}-{\mathcal{K}}^{a}{}_{mcd}{\mathcal{K}}_{an}{}^{cd}-2{\mathcal{K}}^{abc}{}_{m}{\mathcal{K}}_{a}{}_{bcn}\right)$
$\quad\left.+J_{0}\left(8{\mathcal{K}}_{mb}{\mathcal{K}}^{b}_{\ n}-2{\mathcal{K}}{\mathcal{K}}_{mn}\right)\right]~{}.$ (195)
Identifying the term in brackets as $\mathbb{L}_{mn}$ from Eq. (78) results in
$\delta_{g}\mathbb{L}|_{\delta\hat{\Gamma}^{c}_{\ ab}=0}=\delta g^{mn}\mathbb{L}_{mn}~{}.$ (196)
We insert Eqs. (191) and (196) into Eq. (180) and collect all results together into the equation of motion tensor $\mathbb{H}_{mn}$
$\delta_{g}(\sqrt{|g|}\mathcal{L})=$ $\tfrac{\sqrt{|g|}}{2\kappa_{0}}\delta g^{mn}\mathbb{H}_{mn}$ (197)
@@
$0=\mathbb{H}_{mn}=$ $\mathcal{H}_{mn}+\tfrac{2J_{0}\kappa_{0}}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}$ (198)
$+2J_{0}\kappa_{0}\left[{\bar{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}+{\tilde{C}}^{p}_{\ a(m}\tilde{\mathbb{E}}^{*\ \ a}_{n)p}\right]~{}~{}~{}~{}~{}~{}~{}~{}$ (199)
$\mathcal{H}_{mn}=$ $-\mathcal{K}_{mn}+\kappa_{0}\left(2\mathbb{L}_{mn}-g_{mn}\mathcal{L}\right)$ (200)
$-g_{mn}\tfrac{2J_{0}\kappa_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}\left({\mathcal{K}}^{a}{}_{bcd}K^{bcd}\right)$ (201)
@@
To show that $\mathcal{H}_{mn}$ is equivalent to the form in Eq. (76), Eq. (106) is used to relate the ¯${\bar{\nabla}}_{a}$ term to a ˘$\breve{\nabla}_{a}$ term as
¯${\bar{\nabla}}_{a}({\mathcal{K}}^{a}{}_{bcd}K^{bcd})=$ (˘$(\breve{\nabla}_{a}-({\rm d}+1)g_{a})({\mathcal{K}}^{a}{}_{bcd}K^{bcd})$ (202)
$=$ $\tfrac{1}{\sqrt{|g|}}\breve{\nabla}_{a}(\sqrt{|g|}{\mathcal{K}}^{a}{}_{bcd}K^{bcd})$ (203)
where the last equality follows from the definition of $g_{a}$ in Eq. (36) along with the facts that ˘$\breve{\nabla}_{a}$ obeys a Leibniz rule and acts on $|g|$ as merely a partial derivative.
Finally, utilizing Eqs. (54), (136), and (138), we cast Eq. (198) in the form of Einstein equations
$\hat{R}_{mn}-\tfrac{1}{2}g_{mn}\hat{R}=\kappa_{0}\mathbb{T}_{mn}+g_{mn}\Lambda_{0}$ (204)
$\mathbb{T}_{mn}=-\kappa_{0}^{-1}\left[\tilde{Q}_{mn}-\tfrac{1}{2}g_{mn}\tilde{Q}+({\rm d}-1)\left({\bar{{\mathcal{P}}}}_{mn}-\tfrac{1}{2}g_{mn}{\bar{{\mathcal{P}}}}\right)\right]$
$\quad-g_{mn}\tfrac{2J_{0}\lambda_{0}^{2}}{{\rm d}+1}{\bar{\nabla}}_{a}(K_{bcd}\mathcal{K}^{abcd})+2\mathbb{L}_{mn}-g_{mn}\mathbb{L}$ (205)
$+2J_{0}\left[\tfrac{1}{{\rm d}+1}g_{mn}{\bar{\nabla}}_{a}{\bar{\nabla}}_{b}\mathbb{D}^{ab}+{\bar{\nabla}}_{a}\tilde{\mathbb{E}}^{*\ \ a}_{mn}+{\tilde{C}}^{p}_{\ a(m}\tilde{\mathbb{E}}^{*\ \ a}_{n)p}\right].$ (206)
IV Classical Equivalence of Field Equations Across Formulations